OpenStudy (anonymous):

3y - 6/x = 12 find x if y = -4

OpenStudy (anonymous):

3*(-4) -(6/x)=12 -12 - (6/x)=12 -(6/x)=12 +12 -(6/x)=24 6/x = -24 -24x =6 x =-6/24 x = -1/4